OpenStudy (anonymous):

How much power is needed to drive a 1000kg car up a 30 degree incline at a constant speed of 10 m/s? Ignore air resistance and gravity = 10 m/s^2 Answer choices: A) 5000 W B) 10000 W C) 20000 W D) 40000 W

OpenStudy (anonymous):

Power is defined here as\[P = F \cdot v\] The force applied will be equal to the force of gravity on the object (directed along the incline plane). \[F = mg \sin(\theta)\]

OpenStudy (anonymous):

My answer coming 50000 W..